OpenStudy (anonymous):

Simplify: √28/25 A. 4√7 /5 B. √3 C. 4√7 /25<---what i got D. 2√7 /5

OpenStudy (anonymous):

D

OpenStudy (anonymous):

\[\sqrt{\frac{28}{25}}= \frac{\sqrt{28}}{\sqrt{25}}=\frac{\sqrt{2\times 2 \times 7}}{\sqrt{5 \times 5}}\]